How do I figure this out?
2x squared + 5x + 3
I need to make this into a binomial
thank you

Answer by checkley71(8403)

2X^2+5X+3=0
(2X+3)(X+1)=0
2X+3=0
2X=-3
X=-3/2
X=-1.5 ANSWER.
X+1=0
X=-1 ANSWER.